### v4

# Remarks on the Stanley depth and Hilbert depth of monomial ideals with linear quotients

Abstract.
We prove that if $I$ is a monomial ideal with linear quotients in a ring of polynomials $S$ in $n$ indeterminates and $\operatorname{depth}(S/I)=n-2$, then $\operatorname{sdepth}(S/I)=n-2$ and, if $I$ is squarefree, $\operatorname{hdepth}(S/I)=n-2$.
Also, we prove that $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$ for a monomial ideal $I$ with linear quotients which satisfies certain technical conditions.
Key words and phrases:
Stanley depth; Hilbert depth; depth; monomial ideals; linear quotients
2020 Mathematics Subject Classification:
05E40; 06A17; 13A15; 13C15; 13P10
1. Introduction
Let $K$ be a field and let $S=K[x_{1},x_{2},\ldots,x_{n}]$ be the ring of polynomials in $n$ variables. Let $M$ be a $\mathbb{Z}^{n}$-graded $S$-module. A Stanley decomposition of $M$ is a direct sum
$\mathcal{D}:M=\bigoplus_{i=1}^{r}m_{i}K[Z_{i}],$
as $K$-vector spaces, where $m_{i}\in M$, $Z_{i}\subset\{x_{1},\ldots,x_{n}\}$ such that $m_{i}K[Z_{i}]$ is a free $K[Z_{i}]$-module. We define $\operatorname{sdepth}(\mathcal{D})=\min_{i=1}^{r}|Z_{i}|$ and
$\operatorname{sdepth}(M)=\max\{\operatorname{sdepth}(\mathcal{D})|\;\mathcal{D}\text{ is a Stanley decomposition of }M\}.$
The number $\operatorname{sdepth}(M)$ is called the Stanley depth of $M$. Herzog Vlădoiu and Zheng [9] proved that this invariant can be computed in a finite number of steps, when $M=I/J$, where $J\subset I\subset S$ are monomial ideals.
We say that the multigraded module $M$ satisfies the Stanley inequality if
$\operatorname{sdepth}(M)\geq\operatorname{depth}(M).$
[…]
A monomial ideal $I\subset S$ has linear quotients, if there exists $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$, an ordering on the minimal set of generators $G(I)$, such that, for any $2\leq j\leq m$, the ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by variables.
Given a monomial ideal $I\subset S$, Soleyman Jahan [14] noted that $I$ satisfies the Stanley inequality, i.e.
$\operatorname{sdepth}(I)\geq\operatorname{depth}(I).$
However, a similar result for $S/I$, if true, is more difficult to prove, only some particular cases being known. For instance, Seyed Fakhari [8] proved the inequality
$\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$
for weakly polymatroidal ideals $I\subset S$, which are monomial ideals with linear quotients.
In Theorem 2.4, we prove that if $I\subset S$ is a monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$, then $\operatorname{sdepth}(S/I)=n-2$. In Theorem 2.6, we prove that if $I\subset S$ is a monomial ideal with linear quotients which has a Stanley decomposition which satisfies certain conditions, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$. Also, we conjecture that for any monomial ideal $I\subset S$ with linear quotients, there is a variable $x_{i}$ such that $\operatorname{depth}(S/(I,x_{i}))=\operatorname{depth}(S/I)$ and $\operatorname{sdepth}(S/(I,x_{i}))\leq\operatorname{sdepth}(S/I)$. In Theorem 2.12 we prove that if this conjecture is true, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$, for any monomial ideal $I\subset S$ with linear quotients.
Given a finitely graded $S$-module $M$, its Hilbert depth is
$\operatorname{hdepth}(M)=\max\left\{r\;:\;\Large\begin{subarray}{c}\text{ There exists a f.g. graded }S\text{-module }N\\ \text{ with }H_{M}(t)=H_{N}(t)\text{ and }\operatorname{depth}(N)=r\end{subarray}\normalsize\right\}.$
It is well known that $\operatorname{hdepth}(M)\geq\operatorname{sdepth}(M)$. See [3] for further details.
Let $0\subset I\subsetneq J\subset S$ be two squarefree monomial ideals. For any $0\leq j\leq n$, we let $\alpha_{j}(J/I)$ to be the number of squarefree monomials $u\in S$ of degree $j$ such that $u\in J\setminus I$. Also, for $0\leq k\leq d\leq n$, let
$\beta_{k}^{d}(J/I)=\sum_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\alpha_{j}(J/I).$ (1.1)
With the above notations, we proved in [2, Theorem 2.4] that
$\operatorname{hdepth}(J/I)=\max\{d\;:\;\beta_{k}^{d}(J/I)\geq 0\text{ for all }0\leq k\leq d\}.$
Note that, from (1.1), one can easily see that
$\operatorname{hdepth}(J/I)\leq\max\{k\;:\;\alpha_{k}(J/I)>0\}.$ (1.2)
Also, we will make use of the fact that
$\operatorname{hdepth}(J/I)\geq\operatorname{sdepth}(J/I).$ (1.3)
In Section 3 of our paper we study the Hilbert depth of $S/I$, where $I$ is a squarefree monomial ideal with linear quotients. In Proposition 3.2 we compute the numbers $\beta_{k}^{d}(I)$’s and $\beta_{k}^{d}(S/I)$’s. In Corollary 3.3, we express this numbers in terms of hypergeometric functions, thus showing the difficulty in finding explicit formulas for $\operatorname{hdepth}(I)$ and $\operatorname{hdepth}(S/I)$.
The main result of this section is Theorem 3.4, in which we show that if $I$ is a squarefree monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$ then
$\operatorname{hdepth}(S/I)=\operatorname{sdepth}(S/I)=\operatorname{depth}(S/I)=n-2.$
This result makes use of the fact that we can assume $\deg(u_{1})\geq\cdots\geq\deg(u_{m})$; see [15].
2. Main results
Let $I\subset S$ be a monomial ideal and let $G(I)$ be the set of minimal monomial generators of $I$. We recall that $I$ has linear quotients, if there exists a linear order $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$ on $G(I)$, such that for every $2\leq j\leq m$, the ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by a subset of $n_{j}$ variables.
We let $I_{j}:=(u_{1},\ldots,u_{j})$, for $1\leq j\leq m$.
Let $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{j}=\{x_{j}\;|\;x_{j}\notin(I_{j-1}:u_{j})\}$ for $2\leq j\leq m$.
Note that, for any $2\leq j\leq m$, we have
$I_{j}/I_{j-1}=u_{j}(S/(I_{j-1}:u_{j}))=u_{j}K[Z_{j}].$
Hence the ideal $I$ has the Stanley decomposition
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus\cdots\oplus u_{m}K[Z_{m}].$ (2.1)
According to [12, Corollary 2.7], the projective dimension of $S/I$ is
[…]
Hence, Ausländer-Buchsbaum formula implies that
[…]
Note that, (2.1) and (2.2) implies $\operatorname{sdepth}I\geq\operatorname{depth}I$, a fact which was proved in [14]. We recall the following results:
Proposition 2.1.
Let $I\subset S$ be a monomial ideal and $u\in S$ a monomial. Then:
  1. (1)
    $\operatorname{depth}(S/(I:u))\geq\operatorname{depth}(S/I)$. ([11, Corollary 1.3])
  2. (2)
    $\operatorname{sdepth}(S/(I:u))\geq\operatorname{sdepth}(S/I)$. ([5, Proposition 2.7(2)])
Proposition 2.2.
Let $0\to U\to M\to N\to 0$ be a short exact sequence of $\mathbb{Z}^{n}$-graded $S$-modules. Then:
  1. (1)
    $\operatorname{depth}(M)\geq\min\{\operatorname{depth}(U),\operatorname{depth}(N)\}$. (Depth Lemma)
  2. (2)
    $\operatorname{sdepth}(M)\geq\min\{\operatorname{sdepth}(U),\operatorname{sdepth}(N)\}$. ([11, Lemma 2.2])
Proposition 2.3.
Let $I\subset S$ be a proper monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-s-1$, where $10\leq s\leq n-$ 1. Then there exists a subset $\tau\subset[n]$ with $|\tau|=s$ and a monomial $u\in G(I)$, such that $I+(x_{i}\;:\;i\in\tau)=(u)+(x_{i}\;:\;i\in\tau)$.
Proof.
If $I=(u)$ is principal, that is $s=0$, then there is nothing to prove. Assume $s\geq 1$. Since $I$ has linear quotients, we can assume that $G(I)=\{u_{1},\ldots,u_{m}\}$ such that $((u_{1},\ldots,u_{j-1}):u_{j})$ is generated by variables, for every $2\leq j\leq m$. We consider the decomposition (2.1), that is
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus\cdots\oplus u_{m}K[Z_{m}],$
where $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{j}$ is generated by the variables which do not belong to $((u_{1},\ldots,u_{j-1}):u_{j})$, for $2\leq j\leq m$. From (2.2), it follows that $|Z_{j}|\geq s$ for $2\leq j\leq m$. Note that, for any $2\leq j\leq m$, we have
$(u_{1},\ldots,u_{j-1})\cap u_{j}K[Z_{j}]=\{0\}$ (2.3)
We assume, by contradiction, that for any $\tau\subset[n]$ with $|\tau|=s$, there exists $k_{\tau}\neq\ell_{\tau}\in[m]$ such that $u_{k_{\tau}},u_{\ell_{\tau}}\notin(x_{i}\;:\;i\in\tau)$, that is $u_{k_{\tau}},u_{\ell_{\tau}}\in K[x_{i}\;:\;i\notin\tau]$. We claim that there exists $i=i(\tau)\in\tau$ such that $x_{i}\in Z_{k_{\tau}}$ or $x_{i}\in Z_{\ell_{\tau}}$. Indeed, otherwise, since $|Z_{k_{\tau}}|,|Z_{k_{\ell}}|\geq s$ we would have $Z_{k_{\tau}}=Z_{k_{\ell}}=\{x_{i}\;:\;i\notin\tau\}$ and hence
$\operatorname{gcd}(u_{k_{\tau}},u_{\ell_{\tau}})\in u_{k_{\tau}}K[Z_{k_{\tau}}]\cap u_{\ell_{\tau}}K[Z_{\ell_{\tau}}],$
a contradiction. Without any loss of generality, we may assume that $x_{i}\in Z_{k_{\tau}}$.
Let $\sigma=\{n-s+1,\ldots,n\}$. By reordering the variables, we can assume that $k_{\sigma}\geq k_{\tau}$ for any $\tau\subset[n]$ with $|\tau|=s$. Note that $u_{k_{\sigma}}=x_{1}^{a_{1}}\cdots x_{s}^{a_{n-s}}$. By the above argument, there exists $i\geq n-s+1$ such that $x_{i}\in Z_{k_{\sigma}}$. Let $A\subset[n]$ with $|A|=n-s$ such that $i\in A$ and $\{x_{\ell}\;:\;\ell\in A\}\subset Z_{k_{\sigma}}$. It follows that $\tau=[n]\setminus A\neq\sigma$. On the other hand,
$\operatorname{gcd}(u_{k_{\tau}},u_{k_{\sigma}})\in u_{k_{\tau}}S\cap u_{k_{\sigma}}K[Z_{k_{\sigma}}],$
which contradicts (2.3) for $j=k_{\sigma}$. Hence, the proof is complete. ∎
Note that, a monomial ideal $I\subset S$ is principal if and only if $\operatorname{depth}(S/I)=n-1$ if and only if $\operatorname{sdepth}(S/I)=n-1$.
Theorem 2.4.
Let $I\subset S$ be a monomial ideal with linear quotients. If $\operatorname{depth}(S/I)=n-2$, then $\operatorname{sdepth}(S/I)=n-2$.
Proof.
We assume that $G(I)=\{u_{1},\ldots,u_{m}\}$. We use induction on $m$ and $d=\sum_{j=1}^{m}\deg(u_{i})$. If $m=2$, then from [4, Proposition 1.6] it follows that $\operatorname{sdepth}(S/I)=n-2$. If $d=2$, then $I$ is generated by two variables and there is nothing to prove.
Assume $m>2$ and $d>2$. Acording to Proposition 2.3, there exists $i\in[n]$ and $u_{k}\in G(I)$ such that $(I,x_{i})=(u_{k},x_{i})$. Since $(I,x_{i})=(u_{k},x_{i})$, from [4, Proposition 1.2] it follows that $\operatorname{sdepth}(S/(I,x_{i}))\geq n-2$. If $(I:x_{i})$ is principal, then $\operatorname{sdepth}(S/(I:x_{i}))=n-1$. Note that, at least one of $u_{j}$’s must be disivible with $x_{i}$, otherwise we obtain
$G(I,x_{i})=\{u_{1},\ldots,u_{m},x_{i}\},$
a contradiction with the fact that $(I,x_{i})=(u_{k},x_{i})$ and $m\geq 2$. It follows that
$d^{\prime}:=\sum_{u\in G(I:x_{i})}\deg u<d,$
thus, by induction hypothesis, we have $\operatorname{sdepth}(S/(I:x_{i}))=n-2$. In both cases,
$\operatorname{sdepth}(S/(I:x_{i}))\geq n-2.$
From Proposition $1.2(2)$ and the short exact sequence
$0\to S/(I:x_{i})\to S/I\to S/(I,x_{i})\to 0,$
it follows that $\operatorname{sdepth}(S/I)\geq\min\{\operatorname{sdepth}(S/(I:x_{i})),\operatorname{sdepth}(S/(I,x_{i}))\}\geq n-2$. Since $I$ is not principal, it follows that $\operatorname{sdepth}(S/I)=n-2$, as required. ∎
Lemma 2.5.
Let $I\subset S$ be a monomial ideal and $u\in S$ a monomial with $(I:u)=(x_{1},\ldots,x_{m})$. Assume that $S/I$ has a Stanley decomposition
$\mathcal{D}:S/I=\bigoplus_{i=1}^{r}v_{i}K[Z_{i}],$ (2.4)
such that there exists $i_{0}$ with $Z_{i_{0}}=\{x_{m+1},\ldots,x_{n}\}$ and $v_{i_{0}}|u$. Then:
$\operatorname{sdepth}(S/(I,u))\geq\min\{\operatorname{sdepth}(\mathcal{D}),n-m-1\}.$
Proof.
If $\operatorname{sdepth}(S/I)=0$ or $m=n-1$, then there is nothing to prove. We assume that $\operatorname{sdepth}(S/I)\geq 1$ and $m\leq n-2$. Since $S/(I:u)=S/(x_{1},\ldots,x_{m})\cong K[x_{m+1},\ldots,x_{n}]$, from the short exact sequence
$0\to S/(I:u)\stackrel{{\scriptstyle\cdot u}}{{\longrightarrow}}S/I\longrightarrow S/(I,u)\longrightarrow 0,$
it follows that we have the $K$-vector spaces isomorphism
$S/I\cong S/(I,u)\oplus uK[x_{m+1},\ldots,x_{n}].$ (2.5)
From our assumption, $uK[x_{m+1},\ldots,x_{n}]=uK[Z_{i_{0}}]\subset v_{i_{0}}K[Z_{i_{0}}]$. Hence, from (2.4) and (2.5) it follows that
$S/(I,u)\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus\frac{v_{i_{0}}K[Z_{i_{0}}]}{uK[Z_{i_{0}}]}\cong\left(\bigoplus_{i\neq i_{0}}v_{i}K[Z_{i}]\right)\oplus v_{i_{0}}\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]},$ (2.6)
where $w_{0}=\frac{u}{v_{i_{0}}}$. On the other hand, $\operatorname{sdepth}\left(\frac{K[x_{m+1},\ldots,x_{n}]}{w_{0}K[x_{m+1},\ldots,x_{n}]}\right)=n-m-1$. Hence (2.6) yields the required conclusion. ∎
Theorem 2.6.
Let $I\subset S$ be a monomial ideal with linear quotients, $G(I)=\{u_{1},\ldots,u_{m}\}$, $I_{j}=(u_{1},\ldots,u_{j})$ for $1\leq j\leq m$, such that $(I_{j-1}:u_{j})=(\{x_{1},\ldots,x_{n}\}\setminus Z_{j})$, where $Z_{j}\subset\{x_{1},\ldots,x_{n}\}$, for all $2\leq j\leq m$.
We assume that for any $2\leq j\leq m$, there exists a Stanley decomposition $\mathcal{D}_{j-1}$ of $S/I_{j-1}$ such that $\operatorname{sdepth}(\mathcal{D}_{j-1})\geq\operatorname{depth}(S/I_{j-1})$ and there exists a Stanley subspace $w_{j-1}K[W_{j-1}]$ of $\mathcal{D}_{j-1}$ with $w_{j-1}|u_{j}$ and $W_{j-1}=Z_{j}$.
Then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$.
Proof.
From the hypothesis and Lemma 2.5, we have that
$\operatorname{sdepth}(S/I_{j})=\operatorname{sdepth}(S/(I_{j-1},u_{j}))\geq\min\{\operatorname{sdepth}(\mathcal{D}_{j-1}),n-n_{j}-1\}\}\geq$
$\geq\min\{\operatorname{depth}(S/I_{j-1}),n-n_{j}-1\},\text{ for all }2\leq j\leq m,$ (2.7)
where $n_{j}=n-|Z_{j}|$, $1\leq j\leq m$. On the other hand, according to (2.2),
$\operatorname{depth}(S/I)=\min_{j=1}^{m}\{n-n_{j}-1\}.$ (2.8)
Since $\operatorname{sdepth}(S/I_{1})=\operatorname{depth}(S/I_{1})=n-1$, from (2.7) and (2.8) we get the required conclusion. ∎
Example 2.7.
Let $I=(x_{1}^{2},x_{1}x_{2}^{2},x_{1}x_{2}x_{3}^{2})\subset S=K[x_{1},x_{2},x_{3},x_{4}]$. Let $u_{1}=x_{1}^{2}$, $u_{2}=x_{1}x_{2}^{2}$ and $u_{3}=x_{1}x_{2}x_{3}^{2}$. Since $((u_{1}):u_{2})=(x_{1})$ and $((u_{1},u_{2}):u_{3})=(x_{1},x_{2})$, it follows that $I$ has linear quotients with repect to the order $u_{1}\leqslant u_{2}\leqslant u_{3}$. Moreover,
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus u_{3}K[Z_{3}]=x_{1}^{2}K[x_{1},x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]\oplus x_{1}x_{2}x_{3}K[x_{3},x_{4}].$
Let $I_{1}=(u_{1})$ and $I_{2}=(u_{1},u_{2})$. We consider the Stanley decomposition
$\mathcal{D}_{1}:\;S/I_{1}=K[x_{2},x_{3},x_{4}]\oplus x_{1}K[x_{2},x_{3},x_{4}],$
of $S/I_{1}$ with $\operatorname{sdepth}(\mathcal{D}_{1})=\operatorname{sdepth}(S/I_{1})=2$. Let $w_{1}=x_{1}$ and $W_{1}=\{x_{2},x_{3},x_{4}\}$. Clearly, $W_{1}=Z_{2}$ and $w_{1}\mid u_{2}$. As in the proof of Lemma 2.5, we obtain the Stanley decomposition
$\mathcal{D}_{2}:\;S/I_{2}=K[x_{2},x_{3},x_{4}]\oplus\frac{x_{1}K[x_{2},x_{3},x_{4}]}{x_{1}x_{2}^{2}K[x_{2},x_{3},x_{4}]}=K[x_{2},x_{3},x_{4}]\oplus x_{1}K[x_{3},x_{4}]\oplus x_{1}x_{2}K[x_{3},x_{4}].$
Let $w_{2}=x_{1}x_{2}$ and $W_{2}=\{x_{3}\}$. Clearly, $W_{2}=Z_{3}$ and $w_{2}\mid u_{3}$. Hence, according to Theorem 2.6, $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$. In fact, we have that $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)=0$ and
$\mathcal{D}\;:\;S/I=K[x_{2},x_{3},x_{4}]\oplus x_{1}K[x_{3},x_{4}]\oplus x_{1}x_{2}K[x_{4}]\oplus x_{1}x_{2}x_{3}K[x_{4}],$
is a Stanley decomposition of $S/I$ with $\operatorname{sdepth}(\mathcal{D})=\operatorname{sdepth}(S/I)=1$.
We propose the following conjecture:
Conjecture 2.8.
If $I\subset S$ is a proper monomial ideal with linear quotients, then there exists $i\in[n]$ such that $\operatorname{depth}(S/(I,x_{i}))\geq\operatorname{depth}(S/I)$.
Remark 2.9.
Let $I\subset S$ be a monomial ideal with linear quotients, $G(I)=\{u_{1},\ldots,u_{m}\}$, $I_{j}=(u_{1},\ldots,u_{j})$ for $1\leq j\leq m$, such that $(I_{j-1}:u_{j})=(\{x_{1},\ldots,x_{n}\}\setminus Z_{j})$, where $Z_{j}\subset\{x_{1},\ldots,x_{n}\}$, for all $2\leq j\leq m$. $I$ has the Stanley decomposition:
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus\cdots\oplus u_{m}K[Z_{m}],$
where $Z_{1}=\{x_{1},\ldots,x_{n}\}$. We have that
$\operatorname{depth}(S/I)=n-s-1,\text{ where }n-s=\min\{|Z_{j}|\;|\;1\leq j\leq m\}.$
Conjecture 2.8 is equivalent to the fact that there exists $i\in[n]$ such that there is no $1\leq j\leq m$ with $x_{i}\nmid u_{j}$, $x_{i}\in Z_{j}$ and $|Z_{j}|=n-s$.
The following result is well know in literature. However, in order of completeness, we give a proof.
Lemma 2.10.
Let $I\subset S$ be a monomial with linear quotients and $x_{i}$ a variable. Then $(x_{i},I)$ has linear quotients. Moreover, if $S^{\prime}=K[x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]$, then $(x_{i},I)=(x_{i},J)$, where $J\subset S^{\prime}$ is a monomial ideal with linear quotients.
Proof.
We consider the order $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$ on $G(I)$, such that, for every $2\leq j\leq m$, the ideal $(I_{j-1}:u_{j})$ is generated by a nonempty subset $\bar{Z}_{j}$ of variables. We assume that $u_{j_{1}}\leqslant u_{j_{2}}\leqslant\cdots\leqslant u_{j_{p}}$ are the minimal monomial generators of $I$ which are not multiple of $x_{i}$. We have that $((x_{i}):u_{j_{1}})=(x_{i})$. Also, for $2\leq k\leq p$, we claim that
$((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})=(x_{i},\bar{Z}_{j_{k}}).$ (2.9)
Indeed, since $((u_{1},\ldots,u_{j_{k}-1}):u_{j_{k}})=(\bar{Z}_{j_{k}})$ and $x_{i}u_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})$ it follows that $(x_{i},\bar{Z}_{j_{k}})\subset((x_{i},u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})$. Conversely, assume that $v\in S$ is a monomial with $vu_{j_{k}}\in(x_{i},u_{j_{1}},\ldots,u_{j_{k-1}})=(x_{i},u_{1},\ldots,u_{j_{k}-1})$. If $x_{i}\nmid v$, then $vu_{j_{k}}\in(u_{1},\ldots,u_{j_{k}-1})$, hence $v\in(\bar{Z}_{j_{k}})$. If $x_{i}\mid v$, then $v\in(x_{i},u_{1},\ldots,u_{j_{k}-1})$. Hence the claim (2.9) is true and therefore $(x_{i},I)$ has linear quotients. Now, let $J=(u_{j_{1}},\ldots,u_{j_{p}})$. For any $k\geq 2$, we have that
$((u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})\subset((u_{1},\ldots,u_{j_{k}-1}):u_{j_{k}})=(\bar{Z}_{j_{k}}).$ (2.10)
From (2.9) and (2.10), one can easily deduce that $((u_{j_{1}},\ldots,u_{j_{k-1}}):u_{j_{k}})=(\bar{Z}_{j_{k}}\setminus\{x_{i}\})$. Hence, $J$ has linear quotients. ∎
We propose a stronger form of Conjecture 2.8.
Conjecture 2.11.
If $I\subset S$ is a proper monomial ideal with linear quotients, then there exists $i\in[n]$ such that:
  1. i)
    $\operatorname{depth}(S/(I,x_{i}))\geq\operatorname{depth}(S/I)$ and
  2. ii)
    $\operatorname{sdepth}(S/(I,x_{i}))\leq\operatorname{sdepth}(S/I)$.
Note that, if $x_{i}$ is a minimal generator of $I$, then conditions i) and ii) from Conjecture 2.11 are trivial.
Theorem 2.12.
If Conjecture 2.11 is true and $I\subset S$ is a monomial ideal with linear quotients, then $\operatorname{sdepth}(S/I)\geq\operatorname{depth}(S/I)$.
Proof.
We use induction on $n\geq 1$. If $n=1$ then there is nothing to prove. Assume $n\geq 2$. Let $I\subset S$ be a monomial ideal with linear quotients and let $i\in[n]$ such that $\operatorname{depth}(I,x_{i})\geq\operatorname{depth}(I)$ and $\operatorname{sdepth}(S/(x_{i},I))\leq\operatorname{sdepth}(S/I)$. We consider the short exact sequence
$0\to\frac{S}{(I:x_{i})}\to\frac{S}{I}\to\frac{S}{(I,x_{i})}\to 0$ (2.11)
Let $S^{\prime}:=K[x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]$. According to Lemma 2.10, $(x_{i},I)=(x_{i},J)$ where $J\subset S^{\prime}$ is a monomial ideal with linear quotients. Note that:
$\operatorname{sdepth}(S/(x_{i},I))=\operatorname{sdepth}(S/(x_{i},J))=\operatorname{sdepth}(S^{\prime}/J)\text{ and }\operatorname{depth}(S/(I,x_{i}))=\operatorname{depth}(S^{\prime}/J).$
From the induction hypothesis, we have $\operatorname{sdepth}(S^{\prime}/J)\geq\operatorname{depth}(S^{\prime}/J)$. It follows that:
$\operatorname{sdepth}(S/I)\geq\operatorname{sdepth}(S/(I,x_{i}))=\operatorname{sdepth}(S^{\prime}/J)\geq$
$\geq\operatorname{depth}(S^{\prime}/J)=\operatorname{depth}(S/(I,x_{i}))\geq\operatorname{depth}(S/I),$
as required. ∎
Remark 2.13.
Note that, if $I\subset S$ has linear quotients, then $(I:x_{i})$ has not necessarily the same property. For example, the ideal $I=(x_{1}x_{2},x_{2}x_{3}x_{4},x_{3}x_{4}x_{5})\subset K[x_{1},\ldots,x_{5}]$ has linear quotients, but $(I:x_{5})=(x_{1}x_{2},x_{3}x_{4})$ has not. Henceforth, in the proof of Theorem 2.6, we cannot argue, inductively, that $\operatorname{sdepth}(S/(I:x_{i}))\geq\operatorname{depth}(S/(I:x_{i}))$.
3. Remarks on the Hilbert depth
Let $I=(u_{1},\ldots,u_{m})\subset S$ be a squarefree monomial with linear quotients, where $(u_{1},\ldots,u_{i-1}):u_{i}$ is generated by variables for any $2\leq i\leq m$. As we seen in the previous section, $I$ has a decomposition
$I=u_{1}K[Z_{1}]\oplus u_{2}K[Z_{2}]\oplus\cdots\oplus u_{m}K[Z_{m}].$ (3.1)
Moreoever, since $I$ is squarefree, $Z_{1}=\{x_{1},\ldots,x_{n}\}$ and $Z_{i}$ consists in the variables which are not in $(u_{1},\ldots,u_{i-1}):u_{i}$ it follows that $\operatorname{supp}(u_{i})\subset Z_{i}$ for all $1\leq i\leq m$. Therefore, if we denote $d_{i}=\deg(u_{i})$ and $n_{i}=|Z_{i}|$, then $d_{i}\leq n_{i}$, for all $1\leq i\leq n$.
Lemma 3.1.
With the above notations, we have that:
  1. (1)
    $\alpha_{j}(I)=\sum\limits_{i=1}^{m}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq j\leq n$.
  2. (2)
    $\alpha_{j}(S/I)=\binom{n}{j}-\sum\limits_{i=1}^{m}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq j\leq n$.
Proof.
(1) For convenience, we assume that $u_{1}=x_{1}x_{2}\cdots x_{p}$ for some $p\leq n$. For $j\geq p$, a squarefree monomial of degree $j$ in $u_{1}K[Z_{1}]=u_{1}K[x_{1},\ldots,x_{n}]$ is of the form $v=uw$, where $w\in K[x_{p+1},\cdots,x_{n}]$ is squarefree of degree $j-p$. Hence, there are $\binom{n-j}{p-j}=\binom{n_{1}-j}{d_{1}-j}$ such monomials. Similarly, there are $\binom{n_{i}-j}{d_{i}-j}$ squarefree monomials of degree $j$ in $u_{i}K[Z_{i}]$ for all $2\leq i\leq m$. Hence, we get the required conclusion from (3.1).
(2) It follows immediately from (1). ∎
We recall the following combinatorial identity, see for instance [6, Lemma 2.5]:
$\sum_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n}{j}=\binom{n-d+k-1}{k}.$ (3.2)
Now, we state the following result, which follows immediately from Lemma 3.1 and (3.2):
Proposition 3.2.
With the above notations, we have that:
  1. (1)
    $\beta_{k}^{d}(I)=\sum\limits_{i=1}^{r}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq k\leq d\leq n$.
  2. (2)
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}-\sum\limits_{i=1}^{r}\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{n_{i}-j}{d_{i}-j}$ for all $0\leq k\leq d\leq n$.
By straightforward computations, we can deduce the identity
$\sum\limits_{j=0}^{k}(-1)^{k-j}\binom{d-j}{k-j}\binom{N-j}{D-j}=(-1)^{k}\binom{d}{k}\binom{N}{D}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-D}\\ {-d,-N}\end{matrix};{\displaystyle-1}\right],$ (3.3)
where ${}_{r}F_{s}\!\left[\begin{matrix}a_{1},\dots,a_{r}\\ b_{1},\dots,b_{s}\end{matrix};z\right]=\sum\limits_{k=0}^{\infty}\frac{(a_{1})_{k}\cdots(a_{r})_{k}}{k!\,(b_{1})_{k}\cdots(b_{s})_{k}}z^{k}$ is the standard hypergeometric notation; see [13] for futher details.
From Proposition 3.2 and (3.3) we get the following:
Corollary 3.3.
With the above notations, we have that:
  1. (1)
    $\beta_{k}^{d}(I)=(-1)^{k}\binom{d}{k}\sum\limits_{i=1}^{r}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-d_{i}}\\ {-d,-n_{i}}\end{matrix};{\displaystyle-1}\right]$ for all $0\leq k\leq d\leq n$.
  2. (2)
    $\beta_{k}^{d}(S/I)=\binom{n-d+k-1}{k}+(-1)^{k+1}\binom{d}{k}\sum\limits_{i=1}^{r}{}_{3}F_{2}\!\left[\begin{matrix}{1,-k,-d_{i}}\\ {-d,-n_{i}}\end{matrix};{\displaystyle-1}\right]$ for all $0\leq k\leq d\leq n$.
The problem of computing $\operatorname{hdepth}(I)$ and $\operatorname{hdepth}(S/I)$ using directly the formulas given in Corollary 3.3 seems hopeless. However, we can tackle the following particular case:
Theorem 3.4.
Let $I\subset S$ be a squarefree monomial ideal with linear quotients with $\operatorname{depth}(S/I)=n-2$. Then
$\operatorname{hdepth}(S/I)=\operatorname{sdepth}(S/I)=n-2.$
Proof.
From Theorem 2.4 and (1.3) it follows that
$\operatorname{hdepth}(S/I)\geq\operatorname{sdepth}(S/I)=n-2.$
Hence, in order to complete the proof it is enough to show that $\operatorname{hdepth}(S/I)\leq n-2$. If $\alpha_{n-1}(S/I)=0$ then, according to (1.2), there is nothing to prove.
Suppose that $\alpha_{n-1}(S/I)=s>0$. From [15, Lemma 2.1] we can assume that $\deg(u_{1})\geq\deg(u_{2})\geq\cdots\geq\deg(u_{m})$, where $u_{1}\leqslant u_{2}\leqslant\cdots\leqslant u_{m}$ is the linear order on $G(I)$. If $u_{1}=x_{1}x_{2}\cdots x_{n}\in G(I)$ then $I=(u_{1})$ and $\operatorname{depth}(S/I)=n-1$, a contradiction. Therefore:
$\deg(u_{1})=\deg(u_{2})=\cdots=\deg(u_{s})=n-1\text{ and }\deg(u_{i})<n-1\text{ for }s+1\leq i\leq m.$
Let $L=x_{1}x_{2}\cdots x_{n}$. By reordering the variables, we can assume that
$u_{1}=L/x_{n},\;u_{2}=L/x_{1},\;u_{3}=L/x_{2},\ldots,u_{s}=L/x_{s-1}.$
By straightforward computations, it follows that:
$Z_{1}=\{x_{1},\ldots,x_{n}\},\;Z_{2}=Z_{1}\setminus\{x_{1}\},\ldots,\;Z_{s}=Z_{1}\setminus\{x_{s-1}\}.$
We claim that $s=m$. Indeed, if $s<m$ and $\deg(u_{s+1})<n-1$, then the ideal $(u_{1},\ldots,u_{s}):u_{s+1}$ is not generated by variables, a contradiction.
If $s=m=1$ then $I=(u_{1})$ and again we get a contradiction from the fact that $\operatorname{depth}(S/I)=n-1$ in this case. Hence $s\geq 2$ and
$I=(L/x_{n},L/x_{1},\ldots,L/x_{s-1}),\text{ where }L=x_{1}x_{2}\cdots x_{n}.$ (3.4)
From (3.4) and Lemma 3.3 it follows that
$\alpha_{j}(S/I)=\begin{cases}\binom{n}{j},&0\leq j\leq n-2\\ n-s,&j=n-1\\ 0,&j=n\end{cases}.$ (3.5)
From (3.5) we therefore get
$\beta_{n-1}^{n-1}(S/I)=\sum_{j=0}^{n-1}(-1)^{n-1-j}\alpha_{j}(S/I)\leq\sum_{j=0}^{n-1}(-1)^{n-1-j}\binom{n}{j}-2=1-2=-1,$
and therefore $\operatorname{hdepth}(S/I)\leq n-2$, as required. ∎
Remark 3.5.
Note that the inequality $\operatorname{hdepth}(S/I)\geq n-2$ follows also from (3.5), by straightforward computations, so it is not necessarily to use in the proof Theorem 2.4 and the fact that $\operatorname{hdepth}(S/I)\geq\operatorname{sdepth}(S/I)$. Also, from the proof of Theorem 3.4, if $I$ has at least one generator of degree $n-1$ then $I$ is generated in degree $n-1$.
We mention that the squarefree monomial ideals $I$ with linear quotients of degree $n-2$ were extensively studied in [10]. For such an ideal, using a similar argument as in the proof of Theorem 3.4, one can deduce that either $I$ is principal, either $\operatorname{hdepth}(S/I)=n-3$.
Acknowledgements
The second author was supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS - UEFISCDI, project number PN-III-P1-1.1-TE-2021-1633, within PNCDI III.